# A comprehensive review of assistive technologies for children with dyslexia

### v1

@@ 5.2. Tangible User Interface
A group of researchers Jamali et al. introduced the use of tangible user interfaces to enhance learning engagement for children with dyslexia. Tangible objects like cards, toys were integrated with a mobile application, allowing children to interact and practice phonology, spelling, and reading. The engagement and fine motor skills involved in the tangible interaction promote a deeper understanding of sentence structure. The use of TUI with the blocks stimulates hands-on learning and facilitates the development of language skills, including grammar and syntax. The experiment with 30 children using this method resulted higher usability score i.e. 79.5The authors Antle et al. designed a tactile learning system integrating 3D printed tangible letters with dynamic color cues and tactile feedback to support letter recognition and sound association for dyslexic children.The printed letters, have mechanical keys that only particular letters fits on the device slots which helps children with dyslexia to identify mirror words like b and d. Researchers co-designed the activities with teachers to ensure alignment with educational needs and tailored the system to promote accurate letter orientation and sound decoding. While author had presented pilot project the real-world experiment are yet to be conducted however this system in in cooperates innovative approach to multi-sensory learning which helps in active learning among children with dyslexia.
Figure 5. Tangible user interface.
Table 5. Descriptive categorization of tangible User Interface for children with Dyslexia
@@
Technological Feature Tangible User Interface integrated with mobile application for enhanced learning engagement. Text to speech, color codes to identify letters sounds,helps in identifying mirrored letters like (p,q,d,b) to prevent spelling error and provides physical sensation to better understand alphabets.
Targeted Device Mobile tablets. Touch-screen laptop with an input platform and tangible letter blocks.
Targeted Skills Reading, spelling, pronunciation, and writing. Letter recognition, sound decoding, and rectifying mirror letter orientation.
Total Participants 30 children with dyslexia (15 in treatment group, 15 in control group). No participants yet, system co-designed with teachers for future validation.
Validation Methods System Usability Scale (SUS), Again-Again Table, observational notes, and performance checklist. Not specified.
Findings Improved engagement, learning performance, and usability score. Tangible letters with physical constraints may enable children to correctly use canonical orientation for letters, particularly for mirrored ones (p,q,d,b).
Gaming Activities Vowels, syllable, synonyms exercises, Interactive modules for phonology, spelling, and reading with tangible cards and toys. Activities involve arranging 3D printed letters into device slot to form words, with immediate feedback and dynamic color changes.
Software Used Not Specified. Arduino IDE for 3D letter processing and interaction.
5.3. Virtual Reality
@@
A group of researchers Pedroli et al. introduced VR application for rehabilitation, particularly for children with dyslexia. The prime objective of this application was to enhance users engagement and adherence, to address the limitation traditional therapy posses. The application was tailored to make learning effective and enjoyable. Word identification, and retaining focus on children were major activities on this system. In summary significant improvements in word-reading test scores and homophonic writing signifying VR’s potential on rehabilitation an engaging alternative to traditional therapies.
@@
The researchers Saunier et al. experimenting VR technology with Eye-Tracking (ET) and Brain-Computer Interface (BCI), to support children with dyslexia in therapy sessions. This prototype has not be clinically tested however author’s hypothesis on this technology is to monitor attention and cognitive load in real-time, thereby the system dynamically adapting exercises to the learner’s needs, enhancing personalization. This is a pilot study and the research is still on early phase however it can introduce novel method for combining neuroscience and VR to address the challenges faced by children with dyslexia.
@@
A group of researchers Ullah et al., Am Faleel et al. has develop a customizable VR Learning Environment to each basic geometric shapes (cubes, cylinders, pyramids) to children with dyslexia. This application is assists children in in object identification and categorization tasks.Researchers had put forward multi-sensory VR realms which can bolster memory retention and comprehension through simultaneous sensory engagement. Dyslexic children can observe, listen to, and maneuver virtual entities, assisting them in grasping concepts more profoundly and cementing information more efficaciously.
@@
The researchers Maresca et al. emphasizes multisensory learning on immersive VR environments. Researchers implemented user-centered designs to provide interactive reading and phonological tasks, fostering inclusion and motivation in learning processes. A significant difference in the word-reading test scores (p = 0.019) As well as in homophone writing (p = 0.034) was found.This qualitative analyses indicates higher engagement and motivation among users, presenting the effectiveness of multi-sensory VR interventions.
@@
Virtual devices are used by children with dyslexia as a part of rehabilitation therapy. It is particularly used for object identification and adhering attention. The following table reflects the multi-sensory modalities used by virtual reality applications, where it encompasses all four elements.
@@
Figure 6. Multi sensory modalities used on Virtual Reality devices.
@@
Table 6. Descriptive categorization of virtual reality applications for children with dyslexia
Category Pedroli et al. Saunier et al. Ullah et al. Maresca et al.
@@
Research Gap Traditional rehabilitation techniques lack user engagement and adherence due to repetitive tasks. Rehabilitation methods for dyslexia are repetitive and discourage therapy adherence. Limited multisensory techniques to address the challenges with dyslexia. Limited research on VR applications to enhance learning ability.
@@
Technological Feature VR Rehabilitation System with prime focus on cognitive development. Multisensory exercises and biosensor technology,Hybrid VR interface combining eye tracking and brain computer interface. Customizable VR environment to teach object identification and categorization. Immersive and multisensory VR environments with prime focus on reading words.
@@
Total Participants 10 children with dyslexia. Not Specified. Not Specified, Not Specified,designed for children aged 6-12. 28 children(14 control group and 14 experimental group).
Validation Methods Wechsler intelligence scale for children-IV. Pilot test. Qualitative analysis and feedback from participants. Mann–Whitney U-test.
Findings Significant decrease in the time of reading low frequency long words.Virtual Reality devices are efficient in rehabilitation for children with dyslexia on improving attentional skills Multisensory applications can be used for rehabilitation for children with dyslexia.Expected to improve therapy adherence and engagement. Improved engagement, object identification skills, and motivation. Virtual Reality in rehabilitation programs could be one assistive tool for children with dyslexia, classifying it as a promising treatment to enhance language skills, reduce disability, and promote psychological well-being.
Gaming Activities Posner cueing Task, Attentional Blink Task (identifying which letter is white), Reading words and not words(cat and lat). Gamified 3D exercises including to follow object to focus, finding Object in scene, reading exercise. Interactive categorization tasks using shapes and spatial relations. Exercises in which the children interacts with objects, Playful VR environment for motivation,homophone writing exercise.
@@ 5.4. Haptics
The researchers Park et al. introduced haptic technology that aids children with special abilities to improve fine motor control for their handwriting. Multilevel of haptic guidance were provided for various letter depending upon the complexity of alphabets. A group of 42 children participated in this research and the result showed that disturbance haptic guidance was the most effective for high complexity handwriting tasks (such as writing the letters “o” and “g”), partial haptic guidance was the most effective for medium complexity handwriting tasks (such as “t,” “r,” “s,” “e,” “n,” “a,” and “b”), and full haptic guidance was the most effective for low complexity letters (such as “i”). Haptic based technology provides real-time feedback with visual, kinesthetic, and touch sensation which can improve brain activity in learning. The Letters that were frequently used, confusing (b and d) were prioritized to evaluate haptic systems’ effectiveness in enhancing handwriting.
Similarly, Am Faleel et al. introduced Writely, a haptic force feedback system that uses Haptic force feedback for training non-dominant hand writing. Researchers had implemented guidance and anti-guidance system where the children who uses this system would provide a forced haptic feedback if they glides the letters trace. The system is on early stage of exploration, and can establish with a low cost solution for enhancing motor skills for special children.
Researchers Park et al. introduced falcon haptic device used in experiments. Children could grab the pencil-shaped end effector and the haptic device guides the participants’ handwriting to know how to write the task letter or draw task shape. 12 children with cognitive and fine motor delays and mild motor difficulty participated in the study. Haptic based learning guidance was resulted to improve motor functions. Although haptic device were not tailored to children with dyslexia, but all of the technology this technology can be a handy tool for children with dyslexia to enhance their learning abilities. The haptic devices presented all modalities, in enhancing writing abilities. The multi sensory model used on haptic device are shown in table below.
Figure 7. Multi sensory modalities used on haptic devices.
6. Introduction to Non-computing Tools for Dyslexia
@@
The authors Ziadat purposes the VAKT modality which includes Visual, Auditory, kinesthetic and Tactile which provides multi sensory way of learning. The key purpose of the study was to examine how effective is this model in enhancing reading and reading comprehension among student with dyslexia. Several activities included during the experimental session such as word tracing, skywriting, and writing on the wrist. Apart form that word cards, color pens, crayon, mirror, sand, and whiteboard were used to make learning more elegant and interactive. Mirror was used a visual stimulating tool to look to his self when sounded out the phoneme, and simultaneously tapping his fingers to the thumb.39 students participated the survey which resulted statistically significant improvements in comprehension reading of the experimental group referred to the implemented multi-sensory approach (VAKT strategy).
@@
A group of researchers Omar et al. from Malaysia conducted a comparison of visual aids for enhancing reading performance in children with dyslexia.Three specific tools were evaluated: a simple magnifier, which contains a magnifying lens to enlarge text; a visual tracking magnifier, which features a guide line along with a magnifying device that is effective in reducing glare while improving readability; and lastly, a typoscope, which focuses on a line of text by blocking out extra words and glare. 80 school children participated in this case study with critical finding that visual tracking magnifier(VTM) intervention yielded the greatest improvement after 12 weeks of intervention. Hence, author put forward an idea that a VTM can be used as part of a rehabilitation program for all children with dyslexia.
@@
The authors Liu et al. introduced an structural approach to mitigate traditional method of teaching students with dyslexia. The practice of using the whole word to teach reading and writing for children with dyslexia were proven inefficient. The phonics-based program with special including structured, cumulative, multi-sensorial activities stimulate interest, attention as well as cultivate appropriate skills progressively. Several structural activities were designed to acheive the goal. Ranging activities from introduction of simple vowel, consonant letters to smaller phrases and sentence were introduced. Apart from this use of story-based Fitzroy books, laminated abc cards, a word list containing sequence of vowel, consonant alphabets, multi syllables exercise were introduced. Similarly the physical object or materials were used on this process which includes a set of bottle-caps representing each letter (26) in the alphabet in learning alphabets, tactile mat to practice skywriting, white board and whiteboard markers, exercise book, skipping ropes, hula hopes and many more. 28 total students participated on this experiment which resulted as the experiment had improved children progress in reading, spelling, writing and comprehension significantly determined by pre-test and post-test evaluation.
@@
A group of researcher Ferraz et al. conducted an experiment to determine the effects of a phonological remediation on reading and writing program in children with dyslexia. Researchers evaluated phonological cal awareness, through multiple exercise like rapid naming, working memory, reading and writing of words and non words, thematic writing. Each phonological intervention consists of exercise involving phonemes and syllables, syllabic and phonemic manipulation, and rhyme and alliteration. The higher complex exercise includes the addition and subtraction of phonemes and syllables, syllabic and phonemic manipulation, oral comprehension etc.
@@
The authors Heikkila and Knight explored inclusive music teaching strategies to enhance the learning abilities of elementary-aged children with developmental dyslexia. By incorporating rhythmic, melodic, and kinesthetic activities, the study aimed to improve phonological awareness, auditory discrimination, and motor coordination. Techniques included using body movements to represent pitches, clapping rhythms, and transitioning these skills to musical instruments like the keyboard. More descriptive categorization of non computing device is illustrated by the table below.
@@
Table 7. Descriptive categorization of Non-computing devices and it’s qualities for dyslexic students
Non-computing Features Benefits
@@
VAKT Strategy Combines Visual, Auditory, Kinesthetic, and Tactile learning (multisensory approach). Primary activities involved word tracing, skywriting, and writing on the wrist, playing with word cards, color pens, crayon, mirror, sand, and whiteboard for active learning. Improves oral reading and reading comprehension by engaging multiple senses.
@@
Visual Tracking Magnifier (VTM) Magnifies text, provides a clear line guide, and reduces glare with filters. Enhances reading speed and accuracy, reduces visual strain, and helps with line focus.
@@
Simple Magnifier Magnifies text without additional visual aids. Improves readability for children by enlarging text.
@@
Typoscope Frames a specific line or area of text to reduce distractions. Focuses attention on a single line, improving reading comprehension and reducing cognitive load.
@@
SMARTER Phonics Program Structured phonics-based reading program. Included multiple learning activites with the use of set of story-based Fitzroy books,laminated abc card, word list with syllables words, use of for learning alphabets, tactile mat for writing, skipping ropes and hula hoops for interactive learning Enhances decoding skills, phonemic awareness, and confidence in reading and writing, especially for rural and indigenous children.
@@
Phonological Remediation Program Focused on phonological awareness, rapid naming, reading and writing of words and non words, thematic writing. Each phonological intervention consists of exercise involving phonemes, addition and subtraction of phoneme, syllables, syllabic and phone- mic manipulation, and rhyme and alliteration. The higher complex exercise includes the addition and subtraction of phonemes and syllables, syllabic and phonemic manipulation, oral comprehension Shows significant improvement in reading, writing, and memory retention through systematic intervention.
@@
Music Therapy in Education Integrates melodic and rhythmic activities to engage students with activities including , auditory discrimination. Improves auditory processing, phonological skills, and temporal processing abilities, enhancing overall learning engagement.
7. Research Questions

### v2

@@ 5.2 Tangible User Interface
A group of researchers Jamali et al. [31] introduced the use of tangible user interfaces to enhance learning engagement for children with dyslexia. Tangible objects like cards, toys were integrated with a mobile application, allowing children to interact and practice phonology, spelling, and reading. The engagement and fine motor skills involved in the tangible interaction promote a deeper understanding of sentence structure. The use of TUI with the blocks stimulates hands-on learning and facilitates the development of language skills, including grammar and syntax. The experiment with 30 children using this method resulted higher usability score i.e. 79.5%.
The authors Antle et al. [32] designed a tactile learning system integrating 3D printed tangible letters with dynamic color cues and tactile feedback to support letter recognition and sound association for dyslexic children. The printed letters, have mechanical keys that only particular letters fits on the device slots which helps children with dyslexia to identify mirror words like ’b’ and ’d’. Researchers co-designed the activities with teachers to ensure alignment with educational needs and tailored the system to promote accurate letter orientation and sound decoding. While author had presented pilot project the real-world experiment are yet to be conducted however this system in in cooperates innovative approach to multi-sensory learning which helps in active learning among children with dyslexia.
TUI-based interventions are summarized in Figure 5, highlighting the multisensory modalities they employ (visual, auditory, kinesthetic, tactile) and the specific language skills they target and descriptive details in Table 6
Figure 5: Tangible user interface.
Table 6: Descriptive categorization of tangible user interfaces for children with dyslexia.
@@
Technological Features Tangible user interface integrated with a mobile application for enhanced learning engagement. Text-to-speech; colour-coded letter–sound mapping; support for mirrored letters (p, q, d, b) to reduce spelling errors; physical interaction to reinforce letter understanding.
Targeted Device Tablet devices. Touch-screen laptop with an input platform and tangible letter blocks.
Targeted Skills Reading, spelling, pronunciation, and writing. Letter recognition, sound decoding, and rectifying mirror-letter orientation.
Total Participants 30 children with dyslexia (15 treatment, 15 control). No participants yet; system co-designed with teachers for future validation.
Validation Methods System Usability Scale (SUS); Again–Again Table; observational notes; performance checklist. Not specified.
Findings Improved engagement, learning performance, and usability scores. Physical letter constraints may help children use canonical letter orientation, particularly for mirrored letters (p, q, d, b).
Gaming Activities Vowel, syllable, and synonym exercises; interactive modules for phonology, spelling, and reading using tangible cards and toys. Arrange 3D-printed letters into the device slot to form words, with immediate feedback and dynamic colour changes.
Software Used Not specified. Arduino IDE for 3D-letter processing and interaction.
[…]
5.3 Virtual Reality
@@
A group of researchers Pedroli et al. [33] introduced VR application for rehabilitation, particularly for children with dyslexia. The prime objective of this application was to enhance users engagement and adherence, to address the limitation traditional therapy posses. The application was tailored to make learning effective and enjoyable. Word identification, and retaining focus on children were major activities on this system. In summary significant improvements in word-reading test scores and homophonic writing signifying VR’s potential on rehabilitation an engaging alternative to traditional therapies.
@@
The researchers Saunier et al. [34] experimenting VR technology with Eye-Tracking (ET) and Brain-Computer Interface (BCI), to support children with dyslexia in therapy sessions. This prototype has not be clinically tested however author’s hypothesis on this technology is to monitor attention and cognitive load in real-time, thereby the system dynamically adapting exercises to the learner’s needs, enhancing personalization. This is a pilot study and the research is still on early phase however it can introduce novel method for combining neuroscience and VR to address the challenges faced by children with dyslexia.
@@
A group of researchers [35, 36] has develop a customizable VR Learning Environment to each basic geometric shapes (cubes, cylinders, pyramids) to children with dyslexia. This application is assists children in in object identification and categorization tasks.Researchers had put forward multi-sensory VR realms which can bolster memory retention and comprehension through simultaneous sensory engagement. Dyslexic children can observe, listen to, and maneuver virtual entities, assisting them in grasping concepts more profoundly and cementing information more efficaciously.
@@
The researchers Maresca et al. [37] emphasizes multisensory learning on immersive VR environments. Researchers implemented user-centered designs to provide interactive reading and phonological tasks, fostering inclusion and motivation in learning processes. A significant difference in the word-reading test scores (p = 0.019) as well as in homophone writing (p = 0.034) was found. This qualitative analyses indicates higher engagement and motivation among users, presenting the effectiveness of multi-sensory VR interventions.
@@
Virtual devices are used by children with dyslexia as a part of rehabilitation therapy. It is particularly used for object identification and adhering attention. The Figure 6 reflects the multi-sensory modalities used by VR applications, where it encompasses all four elements with details in Table 7
@@
Figure 6: Multi sensory modalities used on Virtual Reality devices.
@@
Table 7: Descriptive categorization of virtual reality applications for children with dyslexia.
Category Pedroli et al. [33] Saunier et al. [34] Ullah et al. [35] Maresca et al. [37]
@@
Research Gap Traditional rehabilitation techniques are repetitive, reducing engagement and adherence. Rehabilitation for dyslexia tends to be repetitive and discourages therapy adherence. Limited multisensory techniques addressing dyslexia-specific challenges. Limited research on VR to enhance learning ability.
@@
Technological Features VR rehabilitation system with a focus on cognitive development. Multisensory exercises with biosensors; hybrid VR interface combining eye tracking and brain–computer interface. Customizable VR environment to teach object identification and categorization. Immersive, multisensory VR environments with a focus on reading words.
@@
Total Participants 10 children with dyslexia. Not specified. Designed for children aged 6–12 (participant count not specified). 28 children (14 control, 14 experimental).
Validation Methods Wechsler Intelligence Scale for Children (WISC–IV). Pilot test. Qualitative analysis and participant feedback. Mann–Whitney U-test.
Findings Significant decrease in time to read low-frequency long words; VR devices improved attentional skills during rehabilitation. Multisensory VR can support rehabilitation for dyslexia; expected to improve adherence and engagement. Improved engagement, object-identification skills, and motivation. VR in rehabilitation shows promise for improving language skills, reducing disability, and promoting psychological well-being.
Gaming Activities Posner cueing task; attentional blink task (identify the white letter); reading words and non-words (e.g., cat/lat). Gamified 3D exercises: following/focusing on moving objects; object-finding; reading exercises. Interactive categorization using shapes and spatial relations. Object interaction in playful VR scenes; homophone writing exercise to maintain motivation.
@@ 5.4 Haptics
The researchers Park et al. [39] introduced haptic technology that aids children with special abilities to improve fine motor control for their handwriting. Multilevel of haptic guidance were provided for various letter depending upon the complexity of alphabets. A group of 42 children participated in this research and the result showed that disturbance haptic guidance was the most effective for high complexity handwriting tasks (such as writing the letters “o” and “g”), partial haptic guidance was the most effective for medium complexity handwriting tasks (such as “t,” “r,” “s,” “e,” “n,” “a,” and “b”), and full haptic guidance was the most effective for low complexity letters (such as “i”). Haptic based technology provides real-time feedback with visual, kinesthetic, and touch sensation which can improve brain activity in learning. The Letters that were frequently used, confusing (b and d) were prioritized to evaluate haptic systems’ effectiveness in enhancing handwriting.
Similarly, Am Faleel et al. [36] introduced Writely, a haptic force feedback system that uses Haptic force feedback for training non-dominant hand writing. Researchers had implemented guidance and anti-guidance system where the children who uses this system would provide a forced haptic feedback if they glides the letters trace. The system is on early stage of exploration, and can establish with a low cost solution for enhancing motor skills for special children.
Researchers Park et al. [8] introduced falcon haptic device used in experiments. Children could grab the pencil-shaped end effector and the haptic device guides the participants’ handwriting to know how to write the task letter or draw task shape. 12 children with cognitive and fine motor delays and mild motor difficulty participated in the study. Haptic based learning guidance was resulted to improve motor functions. Although haptic device were not tailored to children with dyslexia, but all of the technology this technology can be a handy tool for children with dyslexia to enhance their learning abilities. The haptic devices presented all modalities, in enhancing writing abilities. The multi sensory model used on haptic device are shown in Figure 7.
Figure 7: Multi sensory modalities used on haptic devices.
6 Introduction to Non-computing Tools for Dyslexia
@@
The authors Ziadat [40] purposes the VAKT modality which includes Visual, Auditory, kinesthetic and Tactile which provides multi sensory way of learning. The key purpose of the study was to examine how effective is this model in enhancing reading and reading comprehension among student with dyslexia. Several activities included during the experimental session such as word tracing, skywriting, and writing on the wrist. Apart form that word cards, color pens, crayon, mirror, sand, and whiteboard were used to make learning more elegant and interactive. Mirror was used a visual stimulating tool to look to his self when sounded out the phoneme, and simultaneously tapping his fingers to the thumb. Thirty-nine students participated in the survey, which revealed statistically significant improvements in comprehension reading for the experimental group that utilized the implemented multi-sensory approach (VAKT strategy).
@@
A group of researchers Omar et al. [41] from Malaysia conducted a comparison of visual aids for enhancing reading performance in children with dyslexia. Three specific tools were evaluated: a simple magnifier, which contains a magnifying lens to enlarge text; a visual tracking magnifier, which features a guide line along with a magnifying device that is effective in reducing glare while improving readability; and lastly, a typoscope, which focuses on a line of text by blocking out extra words and glare. A case study involving 80 school children revealed that the visual tracking magnifier (VTM) intervention yielded the most significant improvement after 12 weeks of intervention. Hence, author put forward an idea that a VTM can be used as part of a rehabilitation program for all children with dyslexia.
@@
The authors Liu et al. [42] introduced an structural approach to mitigate traditional method of teaching students with dyslexia. The practice of using the whole word to teach reading and writing for children with dyslexia were proven inefficient. The phonics-based program with special including structured, cumulative, multi-sensorial activities stimulate interest, attention as well as cultivate appropriate skills progressively. Several structural activities were designed to acheive the goal. Ranging activities from introduction of simple vowel, consonant letters to smaller phrases and sentence were introduced. Apart from this use of story-based Fitzroy books, laminated abc cards, a word list containing sequence of vowel, consonant alphabets, multi syllables exercise were introduced. Similarly the physical object or materials were used on this process which includes a set of bottle-caps representing each letter (26) in the alphabet in learning alphabets, tactile mat to practice skywriting, white board and whiteboard markers, exercise book, skipping ropes, hula hopes and many more. 28 total students participated on this experiment which resulted as the experiment had improved children progress in reading, spelling, writing and comprehension significantly determined by pre-test and post-test evaluation.
@@
A group of researcher Ferraz et al. [43] conducted an experiment to determine the effects of a phonological remediation on reading and writing program in children with dyslexia. Researchers evaluated phonological awareness, through multiple exercise like rapid naming, working memory, reading and writing of words and non words, thematic writing. Each phonological intervention consists of exercise involving phonemes and syllables, syllabic and phonemic manipulation, and rhyme and alliteration. The higher complex exercise includes the addition and subtraction of phonemes and syllables, syllabic and phonemic manipulation, oral comprehension etc.
@@
The authors Heikkila and Knight [44] explored inclusive music teaching strategies to enhance the learning abilities of elementary-aged children with developmental dyslexia. By incorporating rhythmic, melodic, and kinesthetic activities, the study aimed to improve phonological awareness, auditory discrimination, and motor coordination. Techniques included using body movements to represent pitches, clapping rhythms, and transitioning these skills to musical instruments like the keyboard. More descriptive categorization of non computing device is illustrated by the Table 8.
@@
Table 8: Descriptive categorization of non-computing devices for children with dyslexia.
Non-computing device Features Benefits
@@
VAKT Strategy Multisensory approach combining visual, auditory, kinesthetic, and tactile modalities. Activities include word tracing, skywriting, writing on the wrist, using word cards, color pens/crayons, mirror work, sand writing, and whiteboard practice. Improves oral reading and reading comprehension by engaging multiple senses.
@@
Visual Tracking Magnifier (VTM) Magnifies text; provides a clear line guide; glare reduction via filters. Enhances reading speed and accuracy, reduces visual strain, and helps maintain line focus.
@@
Simple Magnifier Handheld magnifier that enlarges text without additional visual aids. Improves readability by enlarging text for children with visual tracking challenges.
@@
Typoscope Frames a single line or area of text to reduce visual distractions. Focuses attention on one line at a time, improving comprehension and lowering cognitive load.
@@
SMARTER Phonics Program Structured phonics-based reading program. Materials/activities include story-based Fitzroy readers, laminated ABC cards, syllabified word lists, tactile mat for letter formation, and movement-based tasks (e.g., skipping ropes, hula hoops). Enhances decoding skills, phonemic awareness, and confidence in reading/writing—particularly helpful for rural and indigenous learners.
@@
Phonological Remediation Program Targets phonological awareness, rapid naming, and reading/writing of words and non-words. Exercises include phoneme addition/deletion, syllable segmentation/manipulation, rhyme and alliteration; advanced tasks expand these operations and oral comprehension. Shows significant improvement in reading, writing, and memory retention through systematic intervention.
@@
Music Therapy in Education Integrates melodic and rhythmic activities to build auditory discrimination and engagement. Improves auditory processing, phonological skills, and temporal processing, enhancing overall learning engagement.
[…]
7 Research Questions